OpenStudy (anonymous):

A painter leans a 15 ft ladder against a building. The base of the ladder is 6 ft from the building. To the nearest foot, how high on the building does the ladder reach?

OpenStudy (harsimran_hs4):

 use Pythagoras theorem to find x (height) can you proceed now ?

OpenStudy (anonymous):

18?

OpenStudy (harsimran_hs4):

no 15 ^2 = 6^2 + x^2

OpenStudy (anonymous):

9

OpenStudy (harsimran_hs4):

225 = 36 + x^2 225 - 36 = 189 = x^2 x = sqrt(189)

OpenStudy (anonymous):

14

OpenStudy (harsimran_hs4):

13.7

OpenStudy (anonymous):

i know i just rounded it

OpenStudy (anonymous):

thank you

OpenStudy (harsimran_hs4):

great!!